OpenStudy (anonymous):

How to find the value of sin[2cos^-1(-3/5)] without using calculator??

OpenStudy (anonymous):

Let \[\cos^{-1}\frac{-3}{5} = A \implies \cos A = \frac{-3}{5} \implies \cos^2 A = \frac{9}{25}\] \[\cos^2 A + \sin^2 A = 1 \implies \sin A = \sqrt{1-9/25} = \frac{4}{5} \] We cant to find \[\sin 2A \] which is equal to \[2\sin A \cos A \]. Gogogogogogog

OpenStudy (anonymous):

Anwer is -24/25 and when i use with calculator it gives me that number...

OpenStudy (anonymous):

Cool; I did it without a calculator, using the method above, because I'm not a three year old baby.

OpenStudy (anonymous):

So what is your answer at then end of your method?? 4/5??

OpenStudy (anonymous):

ah so just plug 4/5 to the equation i got it thanks

OpenStudy (anonymous):

No.. The answer is 2 * sin A * cos A. And I told you sin A = 4/5 and cos A = 3/5.

OpenStudy (anonymous):

cos A = -3/5 ***

OpenStudy (anonymous):

Thank you